OpenStudy (anonymous):

a 5kg mass resting on a smooth horizontal, frictionless table is connected to a cable which passes over a pulley and then is fastened to a hanging 10kg mass. find the acceleration of the two objects and the tension of the string.

OpenStudy (anonymous):

No, i wont, you cant make me. Im not going to do that.

OpenStudy (anonymous):

Why not?

OpenStudy (anonymous):

T= 5a -T +10g = 10a -15a =10g a=-10g/15 (- a means the larger block is accelerating downwards)

OpenStudy (ujjwal):

 Tension, T=5a=10(g-a) Solve to find a and then T